# Yang-Lee Quantum Criticality in Various Dimensions

### v3

@@ 5.2 Approaching the 3D YL critical point
$r_{T}=\frac{E_{1}^{(\textbf{5})}-E_{0}^{(\textbf{1})}}{E_{0}^{(\textbf{3})}-E_{1}^{(\textbf{1})}}=3,$ (5.8)
The identification of each state is taken from the operator flow from Ising to Yang-Lee, shown in Fig. 8 and Table 9. Notice, that the expression (5.7) does not depend on the scaling dimension of $\phi$, which we will later find. From this, we obtain $ih_{z}^{*}(N)$ for $N=8,9,\dots,14$ and the IR critical point is extrapolated via (3.3) using a $1/N$ fitting. In Fig. 9, we show an example of the thermodynamic critical point extrapolation as a function of $N$ for distinct criticality criteria [^5] For the FSS criterion, we followed [73] and defined $h_{z}^{*}(N)$ as the point where the function $\sqrt{N+1}E_{10}^{(\textbf{1})}(N+1)=\sqrt{N}E_{10}^{(\textbf{1})}(N),$ here we have used that $R\propto\sqrt{N}$ [78]. , we can see they all converge to the same value within $\sim 1\%$, and we notice that $ih_{z}^{\textrm{crit}}$ is far from its pseudo-critical couterpart for small $N$, in contrast to the 2D case, where $ih_{z}^{\textrm{crit}}$ and $ih_{z}^{*}(N)$ lie within a small perturbative regime [76]. We present some of the extrapolated critical points for various $h_{x}$ in Table 10 and we observe that the deeper we go into paramagnetic phase ($h_{x}>3.16$), the better the convergence; however, if we take a very big $h_{x}$ then we start facing numerical convergence problems. Therefore, we chose $h_{x}\in[4,100]$.
$h_{x}$ $4$ $6$ $8$ $10$ $11$
@@
Figure 9: Yang-Lee thermodynamic critical point $h_{z}^{\textrm{crit}}$ extrapolation for $h_{x}=10$ and $V_{0}=4.75$ with $N\in[8,14]$ using different criticality criteria, the extrapolation is done by a linear in $1/N$ fit. The results agree within a $10^{-2}$ error.
@@ 5.3 Extracting the CFT data
The most remarkable feature of the fuzzy sphere model is that it realizes CFTs in radial quantization, where conformal invariant data such as the scaling dimension and OPE coefficients can be directly extracted from simple observables in fuzzy sphere setup. In this subsection we discuss the extraction of these data numerically. Throughout this subsection we pick the critical point of the system at fixed $V_{0}=4.75$, $h_{x}=20$ and tuning $ih_{z}$ to the pseudo critical coupling $ih_{z}^{*}(N)$. The criterion we use for the pseudo critical point is the CFT ratio of the stress tensor (5.7), in other words, we fix $\Delta_{T^{\mu\nu}}=3$. The rest of the CFT data will have finite size deviation from the CFT values, and we study the extrapolation to thermodynamic limit $N\rightarrow\infty$ numerically. We use the public code package (FuzzifiED)[78] and compute the eigensystem through exact diagonalization up to $N=18$. The thermodynamic limit extrapolation is by fitting the data from $N=12$ to $N=18$ as a function of $N$ using models motivated by the finite size scaling properties of CFTs, also known as conformal perturbation theory [79, 17, 18].
CFT systems have many symmetries. In the fuzzy sphere realization this means each conformally invariant quantity has infinitely many ways to be extract. We see remarkable agreement between many distinct ways to evaluate the same quantity, which provide us non-trivial validations of our numeric procedure. In the bulk of this section we will walk through them in details.
@@
Table 11: The CFT data from the fuzzy sphere results. We also list the values from other approaches for comparison. The subscript $[E]$ represent data extracted from eigenvalues. Subscripts $[Z]$ and $[X]$ represent data extracted from the matrix element of the fuzzy sphere $Z_{\ell}$ and $X_{\ell}$ observables, respectively. Other subscripts represent the Padé level used.
In Table 11 we summarize our key numerical results for the CFT data of 3D YL criticality. Some of the quantities have been computed in the literature [41, 45, 46, 54] and we list them in the table for comparison. The data extracted come from (5.15), (5.16), (5.26), (5.31), and (5.39).
5.3.1 Scaling dimension
@@
$S=S_{\rm CFT}+\frac{g_{\phi}}{N^{(\Delta_{\phi}-d)/2}}\int d^{d}x\mathcal{\phi}(x)+\sum_{i}\frac{g_{i}}{N^{(\Delta_{i}-d)/2}}\int d^{d}x{\cal O}_{i}(x)$ (5.10)
where we separate the relevant deformation $\phi$ and other operators ${\cal O}_{i}$ which are all irrelevant. $g_{\phi}$ and $g_{i}$ are dimensionless couplings. Near the criticality, the separation of scales is given by $\frac{\Lambda_{\rm UV}}{\Lambda_{\rm IR}}=R=\sqrt{N}$, which provides the volume dependence of the couplings. The energy eigenvalues of the finite Hamiltonian matches to the CFT Hamiltonian as
$E_{n}=E_{0}+\frac{\nu}{\sqrt{N}}\left(H_{\rm CFT}+\frac{g_{\phi}}{N^{(\Delta_{\phi}-d)/2}}\int d^{d-1}\Omega\,\phi(\Omega)+\sum_{i}\frac{g_{i}}{N^{(\Delta_{i}-d)/2}}\int d^{d-1}\Omega\,{\cal O}_{i}(\Omega)\right)$ (5.11)
where $H_{\rm CFT}|n\rangle=\Delta_{n}$ gives the CFT scaling dimensions. The relevant (irrelevant) couplings $g_{\phi}$ ($g_{i}$) contribute to the leading conformal perturbation corrections to these eigenvalues through diagonal matrix elements with a growing (suppressed) volume dependence.
@@
where $\Delta$ is the scaling dimension prediction extrapolated to the thermodynamic limit. The model is essentially a polynomial in $\frac{1}{R}=\frac{1}{\sqrt{N}}$ except for beginning at power $w$. We choose this model because it takes into account that the leading correction comes in a specific power $N^{(3-\Delta_{{\cal O}_{I}})/2}$ in conformal perturbation theory, and the series expansion after the leading term is agnostic of the precise form of subleading irrelevant deformations and higher order perturbation theory, which we do not know precisely. For the 3D Yang-Lee criticality, it is likely that the irrelevant deformation ${\cal O}_{I}$ is the subleading scalar primary operator, which has a Ginzburg-Landau description $\phi^{3}$. From $(6-\epsilon)$ expansion and from our fuzzy sphere data we see that $\Delta_{\phi^{3}}\approx 4.6$, and we set $w=0.8$. The extrapolations from the fit should not be, and are not, sensitive to the precise value of the input power $w$. For the number of powers we keep, we use $K=3$ which works for most energy levels except for a few. We detect that $K$ is too high if the coefficient of the powers abruptly increases. To validate the fitting scheme, we test it on the difference between two levels $\Delta_{\partial^{2}\phi}-\Delta_{\phi}$ which we expect to be 2. A linear extrapolation is still visibly deviating from $2$. A higer power series fit brings us significantly closer to the theoretical value. The precise value of the power does not matter very much, so the extrapolation from various powers $w$ and truncations $K$ agree to the third digit, though $w=0.8$, the more informed guess from the conformal perturbation theory, is more stable against small change in truncation power.
@@
In Fig. 11 we show the low spectrum of 3D YL criticality as a result of the extrapolation. In the extrapolation scheme (5.14) we treat each eigenvalues independently, so the conformal algebra is not taken as an input except for rescaling the gap between $\Delta_{\phi}$ and $\Delta_{\partial\phi}$ to 1. In the plot we see that relation between $\Delta_{\phi}$ and the descendant levels are precisely integers, making it easy to distinguish primary from descendant levels. The scaling dimension of the stress tensor is fixed at $\Delta_{T_{\mu\nu}}=3$ by the criticality criterion, and all its descendants in this plot are integers to high precision. The plot shows two other primaries, one at $(\ell=0,\Delta\approx 4.6)$ and the other at $(\ell=4,\Delta\approx 4.9)$. The former agrees with the ($6-\epsilon$) expansion prediction of the $\phi^{3}$ operator in the Ginzburg-Landau description, and the latter seems to belong to the leading double trace family $[\phi\phi]_{0,\ell}$ at spin 4, where the stress tensor also reside. We will denote the spin 4 primary as $Q_{\mu\nu\kappa\lambda}$.
@@
In Fig. 12 we take a closer look at the scalar primary operators $\phi$ and $\phi^{3}$. We see that various fit truncation levels agree to high precision. The small fluctuation of the extrapolation with respect to the fitting parameters provide us a heuristic error bar, so we can determine
@@
Figure 13: The difference between several descendant states and their corresponding primary state. The agreement between $\mathop{\rm\Box}\nolimits=\partial_{\mu}\partial^{\mu}$ and $\partial^{2}=(\partial_{\mu}\partial_{\nu}-{\rm trace})$ rely on conformal symmetry. At finite $N$ they receive different finite size correction. We see that the extrapolation to $N\rightarrow\infty$ bring them all to integers.
@@ 5.3.2 OPE coefficients
Apart from the operator scaling dimensions which encodes the conformally invariant 2-point function data, we are also interested in the OPE coefficients which encode the conformally invariant 3-point function data. The most convenient way to extrapolate such information is through state-operator correspondence and study the fuzzy sphere matrix elements of the form
@@
where $|{\cal O},\ell,m\rangle$ is a fuzzy sphere matrix eigenstate in the spin ($\ell,m$) sector and $M_{\ell_{2}m_{2}}$ is a fuzzy sphere observable. In this work we consider the following observables
@@
and we can use it to determine $C_{\phi\phi\phi^{3}}$. The numerical results of these scalar OPE coefficients are shown in Fig. 16. The predictions are
$|C_{\phi\phi\phi}|$ $=1.9696(31)_{[Z]},\quad 1.969(5)_{[X]}$ (5.26)
$|C_{\phi\phi\phi^{3}}|$ $=0.026(7)_{[Z]},\quad 0.0238(15)_{[X]}$
$|C_{\phi\phi^{3}\phi^{3}}|$ $=1.3774(9)_{[Z]},\quad 1.364(7)_{[X]}~{}.$
@@
We conclude this section with an example of scalar-spin-spin 3-point function. In our convention, we normalize a generic $-\ell$ primary to have a standard form 2-point function
@@
The 24-cell or Icositetrachoron is a self-dual regular 4D polytope (polychoron), which has $N=24$ vertices and $E=96$ edges. We can construct this polychoron by defining its vertices in a four-dimensional Cartesian coordinate system. Its $24$ vertices $v_{i}$ (for $i=1,\dots,24$) are obtained by all permutations of coordinates of the vectors (±$(\pm 1,\pm 1,0,0)$. We draw the 24-cell as graph on the $F_{4}$ Coxeter plane in Figure 19.
@@
When the imaginary magnetic field $ih_{z}$ is turned on, we observe a similar behavior for the energy gaps as in the 2D and 3D cases: the two lowest energy levels merge at some value $ih_{z}^{\textrm{merg}}$ We list the values of this merging point in the Table 12.
$h_{x}$ $15$ $20$ $30$ $40$ $50$ $60$ $70$ $80$ $90$ $100$
@@ Appendix A Characters of $W(F_{4})$
We denote representatives of the conjugacy classes of $W(F_{4})$ by $R_{q_{\textrm{L}},q_{\textrm{R}}}$ for proper rotations and by $\tilde{R}_{q_{\textrm{L}},q_{\textrm{R}}}$ for improper rotations [99], where $q_{\textrm{L}}$ and $q_{\textrm{R}}$ are two unit quaternions: $qq^{{\dagger}}=\sum_{i=0}^{3}q_{i}^{2}=1$, where $q=q_{0}+q_{1}\textbf{i}+q_{2}\textbf{j}+q_{3}\textbf{k}$, $q^{{\dagger}}=q_{0}-q_{1}\textbf{i}-q_{2}\textbf{j}-q_{3}\textbf{k}$ and $\textbf{i}^{2}=\textbf{j}^{2}=\textbf{k}^{2}=\textbf{i}\textbf{j}\textbf{k}=-1$. The actions of the proper rotation $R_{q_{\textrm{L}},q_{\textrm{R}}}$ and the improper rotation $\tilde{R}_{q_{\textrm{L}},q_{\textrm{R}}}$ on a vector $q=(q_{0},q_{1},q_{2},q_{3})$ in 4D space, represented as a quaternion, are
@@
Conjugacy class $1$ $4A_{1}$ $2A_{1}$ $A_{2}$ $D_{4}$ $D_{4}(a_{1})$ $\tilde{A}_{2}$ $C_{3}+A_{1}$ $A_{2}+\tilde{A}_{2}$ $F_{4}(a_{1})$ $F_{4}$ $A_{1}+\tilde{A}_{1}$ $B_{2}$ $A_{3}+\tilde{A}_{1}$ $B_{4}$
Size $1$ $1$ $18$ $32$ $32$ $12$ $32$ $32$ $16$ $16$ $96$ $72$ $36$ $36$ $144$

### v4

@@ 5.2 Approaching the 3D YL critical point
$r_{T}=\frac{E_{1}^{(\textbf{5})}-E_{0}^{(\textbf{1})}}{E_{0}^{(\textbf{3})}-E_{1}^{(\textbf{1})}}=3\,.$ (5.8)
The identification of each state is taken from the operator flow from Ising to Yang-Lee, shown in Fig. 8 and Table 9. Note that the expression (5.7) does not depend on the scaling dimension of $\phi$, which we will find later. From this, we obtain $ih_{z}^{*}(N)$ for $N=8,9,\dots,14$ and the IR critical point is extrapolated via (3.3) using $1/N$ linear fit [^6] Although Conformal Perturbation Theory (CPT) predicts the convergence rate of $ih_{z}^{*}(N)$ to $ih_{z}^{\textrm{crit}}$ at large $N$, we observe that for small values of $N$ the points are fitted well by $1/N$ linear fit. . In Fig. 9, we show an example of thermodynamic critical point extrapolation as a function of $N$ for distinct criticality criteria [^7] For the FSS criterion, we followed [80] and defined $h_{z}^{*}(N)$ as the point where $\sqrt{N+1}E_{10}^{(\textbf{1})}(N+1)=\sqrt{N}E_{10}^{(\textbf{1})}(N),$ where we used that $R\propto\sqrt{N}$ [85]. , we can see they all converge to the same value within $\sim 1\%$, and we notice that $ih_{z}^{\textrm{crit}}$ is far from its pseudo-critical couterpart for small $N$, in contrast to the 2D case, where $ih_{z}^{\textrm{crit}}$ and $ih_{z}^{*}(N)$ lie within a small perturbative regime [83]. We present some of the extrapolated critical points for various $h_{x}$ in Table 10 and we observe that the deeper we go into the paramagnetic phase ($h_{x}>3.16$), the better the convergence; however, if we take a very big $h_{x}$ then we start facing numerical convergence problems. Therefore, we chose $h_{x}\in[4,100]$.
$h_{x}$ 44 66 88 $10$ $11$
@@
Figure 9: Yang-Lee thermodynamic critical point $h_{z}^{\textrm{crit}}$ extrapolation for $h_{x}=10$ and $V_{0}=4.75$ with $N\in[8,14]$ using different criticality criteria; the extrapolation is done by a linear fit in $1/N$. The results agree within a $10^{-2}$ error.
@@ 5.3 Extracting the CFT data
The most remarkable feature of the fuzzy sphere model is that it realizes CFTs in radial quantization, where conformal invariant data such as the scaling dimensions and OPE coefficients can be directly extracted from simple observables in the fuzzy sphere setup. In this subsection we discuss the extraction of these data numerically. Throughout this subsection we pick the critical point of the system at fixed $V_{0}=4.75$, $h_{x}=20$ and tuning $ih_{z}$ to the pseudo-critical coupling $ih_{z}^{*}(N)$. The criterion we use for the pseudo-critical point is the CFT ratio of the stress tensor (5.7), in other words, we fix $\Delta_{T^{\mu\nu}}=3$. The rest of the CFT data will have finite size deviation from the CFT values, and we study the extrapolation to thermodynamic limit $N\rightarrow\infty$ numerically. We use a publicly available code package (FuzzifiED)[85] and compute the eigensystem through exact diagonalization up to $N=18$. Extrapolation to the thermodynamic limit is performed by fitting the data from $N=12$ to $N=18$ as a function of $N$ using models motivated by the finite size scaling properties of CFTs, also known as conformal perturbation theory [86, 87, 17, 18].
CFT systems have many symmetries. In the fuzzy sphere realization, this implies that any conformally invariant quantity can be extracted in many ways. We see remarkable agreement between many distinct ways to evaluate the same quantity, which provides nontrivial validation of our numerical procedure. In this section we will walk through these in details.
@@
Table 11: The CFT data from the fuzzy sphere results. We also list the values from other approaches for comparison. The subscript $[E]$ represent data extracted from eigenvalues. Subscripts $[Z]$ and $[X]$ represent data extracted from the matrix element of the fuzzy sphere $Z_{\ell}$ and $X_{\ell}$ observables, respectively. Other subscripts represent the Padé level used. The error bars of the Fuzzy Sphere estimations are drawn from the largest differences of the extrapolations shown in the corresponding plots. They show the scales of possible fitting errors and should not be understood as bounds on the true values.
In Table 11 we summarize our key numerical results for the CFT data of the 3D YL criticality. Some of the quantities have been computed in the literature [41, 45, 46, 57] and we list them in the table for comparison. The data comes from (5.15), (5.16), (5.26), (5.31), and (5.39).
5.3.1 Scaling dimensions
@@
$S=S_{\rm CFT}+\frac{g_{\phi}}{N^{(\Delta_{\phi}-d)/2}}\int d^{d}x\mathcal{\phi}(x)+\sum_{i}\frac{g_{i}}{N^{(\Delta_{i}-d)/2}}\int d^{d}x{\cal O}_{i}(x)\,$, (5.10)
where we separate the relevant deformation $\phi$ and other operators ${\cal O}_{i}$ which are all irrelevant, and $g_{\phi}$ and $g_{i}$ are dimensionless couplings. Near the criticality, the separation of scales is given by $\frac{\Lambda_{\rm UV}}{\Lambda_{\rm IR}}\sim\sqrt{N}$, which provides the volume dependence of the couplings. The energy eigenvalues of the finite Hamiltonian match to the CFT Hamiltonian as
$E_{n}=E_{0}+\frac{\nu}{\sqrt{N}}\left(H_{\rm CFT}+\frac{g_{\phi}}{N^{(\Delta_{\phi}-d)/2}}\int d^{d-1}\Omega\,\phi(\Omega)+\sum_{i}\frac{g_{i}}{N^{(\Delta_{i}-d)/2}}\int d^{d-1}\Omega\,{\cal O}_{i}(\Omega)\right)\,$, (5.11)
where $H_{\rm CFT}|n\rangle=\Delta_{n}|n\rangle$, and $\Delta_{n}$ are the CFT scaling dimensions. The relevant (irrelevant) couplings $g_{\phi}$ ($g_{i}$) contribute to the leading conformal perturbation corrections to these eigenvalues through diagonal matrix elements with a growing (suppressed) volume dependence.
@@
where $\Delta$ is the scaling dimension prediction extrapolated to the thermodynamic limit. The model is essentially a polynomial in $\frac{1}{R}=\frac{1}{\sqrt{N}}$ except for beginning at power $w$. We choose this model because it takes into account that the leading correction comes in a specific power $N^{(3-\Delta_{{\cal O}_{I}})/2}$ in conformal perturbation theory, and the series expansion after the leading term is agnostic of the precise form of subleading irrelevant deformations and higher order perturbation theory, which we do not know precisely. For the 3D Yang-Lee criticality, it is likely that the irrelevant deformation ${\cal O}_{I}$ is the subleading scalar primary operator, which has a Ginzburg-Landau description $\phi^{3}$. From $(6-\epsilon)$ expansion and from our fuzzy sphere data we see that $\Delta_{\phi^{3}}\approx 4.6$, and we set $w=0.8$. The extrapolations from the fit should not be, and are not, sensitive to the precise value of the input power $w$. For the number of powers we keep, we use $K=3$ which works for most energy levels except for a few. We detect that $K$ is too high if the coefficients of the powers abruptly increase. To validate the fitting scheme, we test it on the difference between two levels $\Delta_{\partial^{2}\phi}-\Delta_{\phi}$ which we expect to be 2. A linear extrapolation is still visibly deviating from $2$. A higer power series fit brings us significantly closer to the theoretical value. The precise value of the power does not matter very much, so the extrapolation from various powers $w$ and truncations $K$ agree to the third digit, though $w=0.8$, the more informed guess from the conformal perturbation theory, is more stable against small changes in truncation power.
@@
In Fig. 11 we show the low spectrum of 3D YL criticality as a result of the extrapolation. In the extrapolation scheme (5.14) we treat each eigenvalues independently, so the conformal algebra is not taken as an input except for rescaling the gap between $\Delta_{\phi}$ and $\Delta_{\partial\phi}$ to 1. In the plot we see that relation between $\Delta_{\phi}$ and the descendant levels are precisely integers, making it easy to distinguish primary from descendant levels. The scaling dimension of the stress tensor is fixed at $\Delta_{T_{\mu\nu}}=3$ by the criticality criterion, and all its descendants in this plot are integers to high precision. The plot shows two other primaries, one at $(\ell=0,\Delta\approx 4.6)$ and the other at $(\ell=4,\Delta\approx 4.9)$. The former agrees with the ($6-\epsilon$) expansion prediction of the $\phi^{3}$ operator in the Ginzburg-Landau description, and the latter seems to belong to the leading double trace family $\phi\phi]_{0,\ell}$ at spin 4, where the stress tensor also reside. We denote the spin 4 primary as $Q_{\mu\nu\kappa\lambda}$.
@@
In Fig. 12 we take a closer look at the scalar primary operators $\phi$ and $\phi^{3}$. We see that various fit truncation levels agree to high precision. The small fluctuation of the extrapolation with respect to the fitting parameters provide us with a heuristic error bar, so we can determine
@@
Figure 13: The difference between several descendant states and their corresponding primary state. The agreement between $\mathop{\rm\Box}\nolimits=\partial_{\mu}\partial^{\mu}$ and $\partial^{2}=(\partial_{\mu}\partial_{\nu}-{\rm trace})$ rely on conformal symmetry. At finite NN they receive different finite size correction. We see that the extrapolation to $N\rightarrow\infty$ bring them all to integers.
@@ 5.3.2 OPE coefficients
Apart from the operator scaling dimensions which encode the conformally invariant 2-point function data, we are also interested in the OPE coefficients which encode the conformally invariant 3-point function data. The most convenient way to extrapolate such information is through state-operator correspondence and study the fuzzy sphere matrix elements of the form
@@
where $|{\cal O},\ell,m\rangle$ is a fuzzy sphere matrix eigenstate in the spin ($\ell,m$) sector and $M_{\ell_{2}m_{2}}$ is a fuzzy sphere observable. The Yang-Lee Hamiltonian (5.5) is a complex symmetric operator, therefore the left eigenvectors coincide with the right eigenvectors, $\langle\psi_{i}|=(|\psi_{i}\rangle)^{T}$. With this definition we have $\langle\psi_{i}|\psi_{j}\rangle=\delta_{ij}$. In this work we consider the following observables
@@
and we can use it to determine $C_{\phi\phi\phi^{3}}$. The numerical results of these scalar OPE coefficients are shown in Fig. 16. The predictions are [^8] The error bars are determined by the difference between the linear and quadratic extrapolations. The purpose of these error bars is to show the scale of a possible fitting error, but not to assert that the true value should lie within the error bar. Note that $C_{\phi\phi^{3}\phi^{3}}$ has non-overlapping error bars.
[…]
@@
We conclude this section with an example of scalar-spin-spin 3-point function. In our convention, we normalize a generic spin-$\ell$ primary to have a standard form 2-point function
@@
The 24-cell or Icositetrachoron is a self-dual regular 4D polytope (polychoron), which has $N=24$ vertices and $E=96$ edges. We can construct this polychoron by defining its vertices in a four-dimensional Cartesian coordinate system. Its $24$ vertices $v_{i}$ (for $i=1,\dots,24$) are obtained by all permutations of coordinates of the vectors $(\pm 1,\pm 1,0,0)$. We draw the 24-cell as graph on the $F_{4}$ Coxeter plane in Figure 19.
@@
When the imaginary magnetic field $ih_{z}$ is turned on, we observe a similar behavior for the energy gaps as in the 2D and 3D cases: the two lowest energy levels merge at some value $ih_{z}^{\textrm{merg}}$ . We list the values of this merging point in the Table 12.
$h_{x}$ $15$ 2020 $30$ $40$ $50$ $60$ $70$ $80$ $90$ $100$
@@ Appendix A Characters of $W(F_{4})$
We denote representatives of the conjugacy classes of $W(F_{4})$ by $R_{q_{\textrm{L}},q_{\textrm{R}}}$ for proper rotations and by $\tilde{R}_{q_{\textrm{L}},q_{\textrm{R}}}$ for improper rotations [106], where $q_{\textrm{L}}$ and $q_{\textrm{R}}$ are two unit quaternions: $qq^{{\dagger}}=\sum_{i=0}^{3}q_{i}^{2}=1$, where $q_{0}+q_{1}\textbf{i}+q_{2}\textbf{j}+q_{3}\textbf{k}, q=q$$^{{\dagger}}=q_{0}-q_{1}\textbf{i}-q_{2}\textbf{j}-q_{3}\textbf{k}$ and $\textbf{i}^{2}=\textbf{j}^{2}=\textbf{k}^{2}=\textbf{i}\textbf{j}\textbf{k}=-1$. The actions of the proper rotation $R_{q_{\textrm{L}},q_{\textrm{R}}}$ and the improper rotation $\tilde{R}_{q_{\textrm{L}},q_{\textrm{R}}}$ on a vector $q=(q_{0},q_{1},q_{2},q_{3})$ in 4D space, represented as a quaternion, are
@@
Conjugacy class 11 $4A_{1}$ $_{1} A2A_{2}$ $D_{4}$ $D_{4}(a_{1})$ $\tilde{A}_{2}$ $C_{3}+A_{1}$ $A_{2}+\tilde{A}_{2}$ $F_{4}(a_{1})$ $F_{4}$ $A_{1}+\tilde{A}_{1}$ $B_{2}$ $A_{3}+\tilde{A}_{1}$ $B_{4}$
Size 11 11 $18$ $32$ $32$ $12$ $32$ $32$ $16$ $16$ $96$ $72$ $36$ $36$ $144$